OpenStudy (anonymous):

Which expression shows the result of applying the Commutative Property of Addition to the expression 6 • (1 + 9 + 7 • 2)? 6 • (9 + 7 • 2) 6 • (1 + 7 + 9 • 2) 6 • (1 + 9 + 7) • 2 6 • (9 + 1 + 7 • 2)

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Show your work.

OpenStudy (anonymous):

there's really not much you can show...it's a concept..it's basically asking which equation will give you the same answer as the original question

OpenStudy (angela210793):

6(10+14)=6*24= use calculator